OpenStudy (anonymous):

Find the integrating factor and use it to find the general solution for the following differential equation: y'+1y = -(6+4t)

OpenStudy (loser66):

just apply formula. why is it hard?

OpenStudy (anonymous):

dont know formula

OpenStudy (loser66):

in yours, which is P(t)?

OpenStudy (anonymous):

i know p(t) is 1 and g(t) is -6-4t

OpenStudy (anonymous):

but after that im lost

OpenStudy (loser66):

ok, so \(e^{\int p(t)}\) = ?

OpenStudy (loser66):

take step. \(\huge \int p(t)\) = ??

OpenStudy (loser66):

do it, answer me, just this time, you can do whatever you want

OpenStudy (loser66):

why??? you know p (t ) = 1 and you don't know integral of 1 = ?

OpenStudy (anonymous):

sorry i went away from the pc (t)

OpenStudy (loser66):

I don't understand your answer, write it down completely, please

OpenStudy (anonymous):

∫p(t) = t

OpenStudy (loser66):

so, \[\huge e^{\int p(t)}= e^t, right?\] now, \[\huge\color{red} {e^{-\int (p(t)}}=?\]

OpenStudy (anonymous):

e^t

OpenStudy (loser66):

where is my - sign?

OpenStudy (loser66):

why? don't get what I mean? just put minus sign before t, that's it. Please, be patient. I am serious. I do the step not for fun

OpenStudy (anonymous):

-te^t

OpenStudy (anonymous):

oh ok i didnt realize that was a negative sign

OpenStudy (loser66):